OpenStudy (anonymous):

3 points Save Given the following chemical equation, how many grams of carbon monoxide gas are produced when 50.0 grams of silicon dioxide reacts with 36.0 grams of carbon? SiO2 (s) + 3C (s) SiC (s) + 2CO (g) 56.0 grams 168 grams 46.6 grams 67.3 grams

OpenStudy (anonymous):

the limiting reagent in this reaction is C. The mole ratio of C: CO is 3: 2, thus the moles of C from 36g = 36/12 = 3. moles of CO = 2 mass of CO produced = 2(12+16) = 56g

OpenStudy (anonymous):

So where does 36 come in for the final answer??